# On energy bounds in asymptotically locally AdS spacetimes

### v1

@@ 4.2 Torus
This reproduces the result in [12] - the only difference is the Fefferman-Graham expansion here is Lorentzian, as opposed to the Riemannian asymptotics on $\Sigma_{t}$ required by the initial data point of view adopted by [12].
As mentioned earlier, with a toroidal boundary, negative energy solutions are also possible [27]. These evade the positive energy theorem just proven because the circle has two spin structures - namely periodic and anti-periodic. Constant spinors, $\hat{\psi}_{0}$, are periodic where as the topology of the AdS soliton in [27] admits only the anti-periodic variety. There doesn’t appear to be any spinorial remedy to this issue; however the modified positive energy conjecture of [27] has recently been confirmed by [33] using very different means.
@@ 4.3 Sphere
[…]
@@
[…]
@@
The result produced is formally identical to [12], the only difference being that [12] considered Riemannian asymptotics on an initial data slice. As such, they don’t have any $t$ dependence in their Killing spinors. Adding this dependence was essentially the main result of [34]. Theorem 4.4 differs from their work in that the $t$ dependence has been extracted from the physical quantities; in contrast, their analogues of $P_{I}$ and $K_{I}$ have explicit $t$ dependence in their definition. The presentation in theorem 4.4 also justifies some claims which were already used in [35].
The eigenvalues of the matrix in theorem 4.4 can’t be found analytically in general and thus there is no concrete inequality, like in theorem 4.2 for example. However, more progress can be made in specific examples. For example, if $n=4$ and $K_{I}=P_{I}=0$, then [^12] Note that it is always possible to choose a frame in which $P_{I}=0$ by performing a conformal isometry of the round sphere boundary [12]. the eigenvalues of $EI+\frac{\mathrm{i}}{2}J_{IJ}\gamma^{0}\gamma^{IJ}$ are $E\pm\sqrt{\frac{1}{2}J_{IJ}J^{IJ}}=E\pm|J|$, leading to the familiar 4D BPS inequality,
@@
[…]
@@
However, the individual terms in the integrand can be analysed further. When $h=s$, the cross-terms between $\hat{\varepsilon}_{h}^{(+)}$ and $\hat{\varepsilon}_{h}^{(-)}$ are no longer zero. The $p_{0}(\hat{\varepsilon}_{h}^{(+)\dagger}\hat{\varepsilon}_{h}^{(+)}+\hat{\varepsilon}_{h}^{(-)\dagger}\hat{\varepsilon}_{h}^{(-)})$ terms produce energy as before, but this time the $p_{0}(\mathrm{ie}^{\mathrm{i}t}\hat{\varepsilon}_{h}^{(-)\dagger}\hat{\varepsilon}_{h}^{(+)}-\mathrm{ie}^{-\mathrm{i}t}\hat{\varepsilon}_{h}^{(+)\dagger}\hat{\varepsilon}_{h}^{(-)})$ terms produce linear momentum, the $\mathrm{i}p_{A}(\hat{\varepsilon}_{h}^{(+)\dagger}\hat{\varepsilon}_{h}^{(+)}+\hat{\varepsilon}_{h}^{(-)\dagger}\hat{\varepsilon}_{h}^{(-)})$ terms produce angular momentum and the $p_{A}(\mathrm{e}^{\mathrm{i}t}\hat{\varepsilon}_{h}^{(-)\dagger}\hat{\gamma}^{A}\hat{\varepsilon}_{h}^{(+)}-\mathrm{e}^{-\mathrm{i}t}\hat{\varepsilon}_{h}^{(+)\dagger}\hat{\gamma}^{A}\hat{\varepsilon}_{h}^{(-)})$ terms produce the centre of mass position. This approach with $\hat{\varepsilon}_{h}^{(\pm)}$ wasn’t followed earlier for computational ease, but also because the interpretation of the physical quantities is much easier in the approach taken earlier.
@@ 4.4 $L(p,1)$
One could consider $L(p,q)$ more generally. However, a similar procedure to the previous lemma shows non-trivial Killing spinors only exist for $q=\pm 1$. L(p,1)L(p,-1) differs from $L(p,1)$ only by a discrete isometry though. For completeness, it should also be noted that $L(p,1)$ admits a unique spin structure for odd $p$ and two inequivalent spin structures for even $p$ [37].
@@
It happens that there is an explicitly known soliton solution with $L(p,1)$ asymptotics [38]. However, it has negative energy, in violation of the theorem just proven. The situation is analogous to the AdS soliton, in that depending on $p$ the solution in [38] either has no spin structure or a different spin structure to the one required by theorem 4.8.
@@ 5 BPS inequalities
[…]
@@
[…]
@@
The only previous work on classical BPS inequalities in these two gauged supergravity theories appears to be in [19, 20, 21]. [19] doesn’t consider magnetic fields in their positive energy theorem at all, so the issues discussed don’t arise. Meanwhile, the result found here disagrees with [20], who explicitly have magnetic charge in their equation 27. However, as explained in [39], this is likely due to some error relating to the fact their analogue of $\nabla_{a}$ omits the $\mathrm{i}A_{a}I$ term and is therefore not gauge covariant. Furthermore, the present results are consistent with subsequent work in [42, 43], where the BPS limit is explicitly required to have zero magnetic charge. Another subtlety explained by [43] is that the 4D minimal supergravity has two distinct vacuum states, one of which has a non-zero magnetic charge determined by the cosmological constant. The connection in [20] is studied in much more detail by [21] and appears to also get different results. Since $\mathrm{i}A_{a}I$ is omitted, the analogue of $\mathbb{M}$ found in [21] is only non-negative definite when a modified dominant energy condition holds - their equation 1.2. However, their condition is very unnatural; for example, it can be seen from the constraint equations that in a purely electrovacuum spacetime with non-zero magnetic field, their modified dominant energy condition never holds.
@@
It turns out an explicit locally supersymmetric solution - see [48] or appendix B of [49] - is known to 5D minimal gauged supergravity with $\mathbb{R}\times L(p,1)$ conformal infinity. However, that solution satisfies a different BPS-like equation, namely
@@
Finally, the only other relevant boundary considered in section 4 is the torus. In this case, theorem 4.2 is unchanged by the electromagnetic fields in either theory because $\bar{\varepsilon}_{k}\varepsilon_{k}$ is simply zero, for the required $\varepsilon_{k}$.
@@ 6 Acknowledgements
Most of all, I would like to thank my supervisor, James Lucietti, for guidance and many discussions throughout this project. I would also like to thank Hari Kunduri, Harvey Reall and especially Piotr Chrusciel for comments and discussions based on earlier drafts of this paper. Finally, I would like to thank the University of Edinburgh’s School of Mathematics for my PhD studentship funding.
@@ Appendix A Inverting the Dirac operator
The key technical requirement is a weighted Poincaré inequality. From section 9 of [51], for the chosen asymptotics, one can develop an inequality,
@@
for any $\psi\in C_{c}^{\infty}$ and some $L^{1}_{\mathrm{loc}}$ function, $w$. Thus,
@@
Suppose $\exists\,\tilde{\mathcal{A}}_{a}$ such that $(\gamma^{I}\mathcal{A}_{I})^{\dagger}=\gamma^{I}\tilde{\mathcal{A}}_{I}$, $\gamma^{IJ}\tilde{\mathcal{A}}_{J}$ is hermitian and
is non-negative definite. Then, $\mathfrak{D}$ is a continuous, linear isomorphism between $\mathcal{H}$ and $L^{2}$.
@@
Define a new connection on spinors, $\widetilde{\nabla}_{a}$, by $\widetilde{\nabla}_{a}=D_{a}-\frac{\mathrm{i}}{2}\gamma_{a}+\tilde{\mathcal{A}}_{a}$ and a new Dirac operator, $\widetilde{\mathfrak{D}}=\gamma^{I}\widetilde{\nabla}_{I}$. Then, equation A.8 can be re-written as
@@
Since $\psi$ could be any compactly supported spinor, it follows that $\Phi$ is a weak solution to $\widetilde{\mathfrak{D}}(\Phi)=0$. Thus, the surjectivity proof reduces to showing $\widetilde{\mathfrak{D}}$ has trivial kernel. Continuing with the new connection, following the same steps as theorem 2.6, one finds if $\psi\in C_{c}^{\infty}$, then
@@
As $\widetilde{\mathbb{M}}$ is assumed to be non-negative definite, this is formally identical to the original Lichnerowicz identity. Then, one can proceed analogously to the analysis of $\mathfrak{D}$ itself by defining an inner product analogous to equation A.1 and thereby concluding that $\widetilde{\mathfrak{D}}$ is injective. Hence, $\widetilde{\mathfrak{D}}(\Phi)=0$ implies $\Phi=\theta-\mathfrak{D}(\varphi)=0$ and therefore $\mathfrak{D}$ is surjective. ∎
At a high level, the previous theorem shows that $\mathfrak{D}$’s injectivity is straightforward, while its surjectivity is equivalent to showing the adjoint, $\mathfrak{D}^{\dagger}$, is injective. The assumptions of the previous theorem are one example which lead to $\mathfrak{D}^{\dagger}$ taking a particularly simple form and having an injectivity proof relying very simply on methods already developed. However, there may be choices of $\mathcal{A}_{a}$ for which no simple $\tilde{\mathcal{A}}_{a}$ exists. Furthermore, even if $\tilde{\mathcal{A}}_{a}$ exists, it may be that $\widetilde{\mathbb{M}}$ is not non-negative definite.
@@
The situation can be rectified if further assumptions are made. The simplest, but most unsatisfying, assumption would be to restrict to electromagnetic fields which have vanishing Poynting vector.

### v2

@@ 4.2 Torus
This reproduces equation 4.14 of [12] - the only difference is the Fefferman-Graham expansion here is Lorentzian, as opposed to the Riemannian asymptotics on $\Sigma_{t}$ required by the initial data point of view adopted by [12].
@@ 4.3 Sphere
$K_{I}$ $=\frac{n-1}{16\pi}\int_{S^{n-2}}f_{(n-1)0\alpha}\frac{\partial\theta^{\alpha}}{\partial x^{J}}\bigg|_{\rho=1}\left(\delta^{{{J}\mathchoice{\makebox[4.08821pt][c]{$\displaystyle$}}{\makebox[4.08821pt][c]{$\textstyle$}}{\makebox[2.53847pt][c]{$\scriptstyle$}}{\makebox[1.81319pt][c]{$\scriptscriptstyle$}}}}_{{\mathchoice{\makebox[5.08472pt][c]{$\displaystyle$}}{\makebox[5.08472pt][c]{$\textstyle$}}{\makebox[3.1884pt][c]{$\scriptstyle$}}{\makebox[2.27742pt][c]{$\scriptscriptstyle$}}{I}}}-\hat{x}^{J}\hat{x}_{I}\right)\mathrm{d}A(s).$ (4.18)
@@
Theorem 2.9 was derived in a vielbein adapted to the Fefferman-Graham expansion. However, since $\bar{\varepsilon}_{k}\gamma^{a}\varepsilon_{k}$ in theorem 2.9 transforms as a Lorentz vector, it suffices to apply $\bar{\varepsilon}_{k}\gamma^{a}\varepsilon_{k}=\Lambda^{{{a}\mathchoice{\makebox[3.51666pt][c]{$\displaystyle$}}{\makebox[3.51666pt][c]{$\textstyle$}}{\makebox[2.1029pt][c]{$\scriptstyle$}}{\makebox[1.50208pt][c]{$\scriptscriptstyle$}}}}_{{\mathchoice{\makebox[4.33765pt][c]{$\displaystyle$}}{\makebox[4.33765pt][c]{$\textstyle$}}{\makebox[2.59009pt][c]{$\scriptstyle$}}{\makebox[1.85005pt][c]{$\scriptscriptstyle$}}{b}}}\bar{\varepsilon}_{k}^{\prime}\gamma^{b}\varepsilon_{k}^{\prime}$. From equation 4.11, one finds
@@
The result produced is formally identical to [12] - in particular, see the unnamed equation directly above their equation 3.12 on their page 11 - the only difference being that [12] considered Riemannian asymptotics on an initial data slice. As such, they don’t have any $t$ dependence in their Killing spinors. Adding this dependence was essentially the main result of [39]. Theorem 4.4 differs from their work in that the $t$ dependence has been extracted from the physical quantities; in contrast, their analogues of $P_{I}$ and $K_{I}$ have explicit $t$ dependence in their definition. The presentation in theorem 4.4 also justifies some claims which were already used in section 6 of [40].
The eigenvalues of the matrix in theorem 4.4 can’t be found analytically in general and thus there is no concrete inequality, like in theorem 4.2 for example. However, more progress can be made in specific examples. For example, if $n=4$ and $K_{I}=P_{I}=0$, then [^15] As explained later, it is typically possible to choose a frame in which $P_{I}=0$ by performing a conformal isometry of the round sphere boundary [12]. the eigenvalues of $EI+\frac{\mathrm{i}}{2}J_{IJ}\gamma^{0}\gamma^{IJ}$ are $E\pm\sqrt{\frac{1}{2}J_{IJ}J^{IJ}}=E\pm|J|$, leading to the familiar 4D BPS inequality,
@@
$\implies\varepsilon_{k}$ $=\frac{1}{\sqrt{1-\rho^{2}}}\left(I-\mathrm{i}\Lambda^{{{I}\mathchoice{\makebox[5.08472pt][c]{$\displaystyle$}}{\makebox[5.08472pt][c]{$\textstyle$}}{\makebox[3.1884pt][c]{$\scriptstyle$}}{\makebox[2.27742pt][c]{$\scriptscriptstyle$}}}}_{{\mathchoice{\makebox[4.08821pt][c]{$\displaystyle$}}{\makebox[4.08821pt][c]{$\textstyle$}}{\makebox[2.53847pt][c]{$\scriptstyle$}}{\makebox[1.81319pt][c]{$\scriptscriptstyle$}}{J}}}x^{J}\gamma_{I}\right)\mathrm{e}^{\mathrm{i}\gamma^{0}t/2}\mathrm{e}^{\theta_{2}\gamma^{2}\gamma^{1}/2}\cdots\mathrm{e}^{\theta_{n-1}\gamma^{n-1}\gamma^{n-2}/2}\varepsilon_{0}.$ (4.36)
@@
However, the individual terms in the integrand can be analysed further. When $h=s$, the cross-terms between $\hat{\varepsilon}_{h}^{(+)}$ and $\hat{\varepsilon}_{h}^{(-)}$ are no longer zero because the Obata theorem no longer applies. The $p_{0}(\hat{\varepsilon}_{h}^{(+)\dagger}\hat{\varepsilon}_{h}^{(+)}+\hat{\varepsilon}_{h}^{(-)\dagger}\hat{\varepsilon}_{h}^{(-)})$ terms produce energy as before, but this time the $p_{0}(\mathrm{ie}^{\mathrm{i}t}\hat{\varepsilon}_{h}^{(-)\dagger}\hat{\varepsilon}_{h}^{(+)}-\mathrm{ie}^{-\mathrm{i}t}\hat{\varepsilon}_{h}^{(+)\dagger}\hat{\varepsilon}_{h}^{(-)})$ terms produce linear momentum, the $\mathrm{i}p_{A}(\hat{\varepsilon}_{h}^{(+)\dagger}\hat{\varepsilon}_{h}^{(+)}+\hat{\varepsilon}_{h}^{(-)\dagger}\hat{\varepsilon}_{h}^{(-)})$ terms produce angular momentum and the $p_{A}(\mathrm{e}^{\mathrm{i}t}\hat{\varepsilon}_{h}^{(-)\dagger}\hat{\gamma}^{A}\hat{\varepsilon}_{h}^{(+)}-\mathrm{e}^{-\mathrm{i}t}\hat{\varepsilon}_{h}^{(+)\dagger}\hat{\gamma}^{A}\hat{\varepsilon}_{h}^{(-)})$ terms produce the centre of mass position. This approach with $\hat{\varepsilon}_{h}^{(\pm)}$ wasn’t followed earlier for computational ease, but also because the interpretation of the physical quantities is much easier in the approach taken earlier.
Finally, consider asymptotic symmetries for these metrics, following the discussion at the end of section 3. Again, the Obata theorem doesn’t hold. Therefore, this time equation 3.58 no longer implies $\xi^{1}_{(0)}=0$. Consequently, equation 3.57 implies $\xi^{\alpha}_{(0)}\partial_{\alpha}$ is a conformal Killing vector of the sphere with $\hat{D}_{\alpha}^{(h)}\xi^{\alpha}_{(0)}=-(n-2)\xi^{1}_{(0)}$. The net effect is the asymptotic symmetry group is equal to the conformal group of $S^{(n-2)}$, namely O(n1,1)O(n-1,1). This reproduces the result of [11]. Furthermore, [11] goes on to show that under these asymptotic symmetries, $E$ & $P_{I}$ can be combined into an object, $P_{a}\equiv(E,P_{I})$, transforming as a vector under O(n1,1)O(n-1,1) and likewise $K_{I}$ & $J_{IJ}$ can be combined into an object, $J_{ab}$ with $J_{0I}=K_{I}$, transforming as a 2-form under O(n1,1)O(n-1,1). Therefore, $E$, $P_{I}$, $J_{IJ}$ and $K_{I}$ are not individually geometric invariants when the cross-section is a round sphere - for example, if $P_{a}$ is timelike, one could always perform an O(n1,1)O(n-1,1) transformation to set $P_{I}$ to zero. As explained in [11, 12], this is fine because various different combinations of $E$, $P_{I}$, $J_{IJ}$ and $K_{I}$ are invariants, e.g. $E^{2}-P_{I}P^{I}$, and theorem 4.4 remains true despite the action of asymptotic symmetries.
@@ 4.4 $L(p,1)$
One could consider $L(p,q)$ more generally. However, a similar procedure to the previous lemma shows non-trivial Killing spinors only exist for $q=\pm 1$. L(p,1)L(p,-1) differs from $L(p,1)$ only by a discrete isometry though.
@@
4.5 Compatibility of spin structures
The entire formalism explored in this work relies on the existence of background Killing spinors near $\mathcal{I}$. By construction, all the examples in this section satisfy that requirement.
However, there are more subtle issues which may arise. Let $\mkern 1.8mu\overline{\mkern-1.8muM\mkern-1.8mu}\mkern 1.8mu$ be an open neighbourhood of $\mathcal{I}$ and let $\bar{g}$ be the background metric. The main problem is that $(\mkern 1.8mu\overline{\mkern-1.8muM\mkern-1.8mu}\mkern 1.8mu,\bar{g})$ may admit multiple spin structures and the spin structure which admits a non-zero solution, $\varepsilon_{k}$, may not be compatible with the spin structure on $(M,g)$. The classic example of this is the AdS soliton [42], which is a vacuum solution of the Einstein equation with $\Lambda<0$ and has metric,
$g$ $=-r^{2}\mathrm{d}t\otimes\mathrm{d}t+\frac{\mathrm{d}r\otimes\mathrm{d}r}{r^{2}(1-1/r^{n-1})}$
$\,\,\,\,\,\,\,+r^{2}\left(\left(1-\frac{1}{r^{n-1}}\right)\mathrm{d}\phi_{2}\otimes\mathrm{d}\phi_{2}+\mathrm{d}\phi_{3}\otimes\mathrm{d}\phi_{3}+\cdots\mathrm{d}\phi_{n-1}\otimes\mathrm{d}\phi_{n-1}\right).$ (4.75)
The angles are identified by $\phi_{2}\sim\phi_{2}+\frac{4\pi}{n-1}$ and $\phi_{3}\sim\phi_{3}+a_{3},\,\cdots,\phi_{n-1}\sim\phi_{n-1}+a_{n-1}$ for arbitrary $a_{3},\,\cdots,a_{n-1}$. The manifold is therefore asymptotically AdS with toroidal cross-section. However, its energy is
$E$ $=-\frac{1}{4}a_{3}\cdots a_{n-1}<0,$ (4.76)
seemingly contradicting theorem 4.2. The reason for this is that a circle admits two inequivalent spin structures - periodic and anti-periodic. Thus, $(\mkern 1.8mu\overline{\mkern-1.8muM\mkern-1.8mu}\mkern 1.8mu,\bar{g})$ admits $2^{n-2}$ inequivalent spin structures. However, the constant spinor, $\hat{\psi}_{0}$, used in the proof of theorem 4.2 is manifestly periodic around each circle of $T^{n-2}$. Therefore, for theorem 4.2 to apply, this particular spin structure must extend beyond $(\mkern 1.8mu\overline{\mkern-1.8muM\mkern-1.8mu}\mkern 1.8mu,\bar{g})$ to all of $(M,g)$. It turns out the global topology of the AdS soliton, namely $\mathbb{R}^{3}\times T^{n-3}$, requires anti-periodicity around $\phi_{2}$. Hence, the AdS soliton’s spin structure is incompatible with the spin structure required for theorem 4.2. There doesn’t appear to be any spinorial remedy to this issue; however the modified positive energy conjecture of [42] has recently been confirmed by [43] using very different means.
Similar issues can arise for the lens space cross-section used in theorem 4.8. Firstly, $L(p,1)$ admits a unique spin structure for odd $p$ and two inequivalent spin structures for even $p$ [44]. It happens that there is an explicitly known soliton solution with $L(p,1)$ asymptotics [45]. However, it has negative energy, in violation of theorem 4.8. The situation is similar to the AdS soliton. This time, depending on $p$ the solution in [45] either has no spin structure or a different spin structure to the one required by theorem 4.8.
@@ 5 BPS inequalities
$R_{ab}-\frac{1}{2}Rg_{ab}+\Lambda g_{ab}$ $=2\bigg(F^{{\mathchoice{\makebox[4.33765pt][c]{$\displaystyle$}}{\makebox[4.33765pt][c]{$\textstyle$}}{\makebox[2.59009pt][c]{$\scriptstyle$}}{\makebox[1.85005pt][c]{$\scriptscriptstyle$}}{c}}}_{{{a}\mathchoice{\makebox[3.57375pt][c]{$\displaystyle$}}{\makebox[3.57375pt][c]{$\textstyle$}}{\makebox[2.1205pt][c]{$\scriptstyle$}}{\makebox[1.51463pt][c]{$\scriptscriptstyle$}}}}F_{bc}-\frac{1}{4}g_{ab}F^{cd}F_{cd}\bigg),$ (5.3)
@@
$R_{ab}-\frac{1}{2}Rg_{ab}+\Lambda g_{ab}$ $=2\bigg(F^{{\mathchoice{\makebox[4.33765pt][c]{$\displaystyle$}}{\makebox[4.33765pt][c]{$\textstyle$}}{\makebox[2.59009pt][c]{$\scriptstyle$}}{\makebox[1.85005pt][c]{$\scriptscriptstyle$}}{c}}}_{{{a}\mathchoice{\makebox[3.57375pt][c]{$\displaystyle$}}{\makebox[3.57375pt][c]{$\textstyle$}}{\makebox[2.1205pt][c]{$\scriptstyle$}}{\makebox[1.51463pt][c]{$\scriptscriptstyle$}}}}F_{bc}-\frac{1}{4}g_{ab}F^{cd}F_{cd}\bigg),$ (5.5)
@@
The only previous work on classical BPS inequalities in these two gauged supergravity theories appears to be in [19, 20, 21, 22]. [19] doesn’t consider magnetic fields in their positive energy theorem at all, so the issues discussed don’t arise.
Meanwhile, the result found here disagrees with [20], who explicitly have magnetic charge in their equation 27. However, as explained in [46], this is likely due to some error relating to the fact their analogue of $\nabla_{a}$ - see their equation 24 - omits the $\mathrm{i}A_{a}I$ term and is therefore not gauge covariant. Furthermore, the present results are consistent with subsequent work in [49, 50], where the BPS limit is explicitly required to have zero magnetic charge. Another subtlety explained by [50] is that the 4D minimal supergravity has two distinct vacuum states, one of which has a non-zero magnetic charge determined by the cosmological constant.
[21] uses the same connection as [20] and studies it in much greater detail. Despite using the same connection, their main result - their theorem 1.1 - differs from the main result of [20] - their equation 27. Since $\mathrm{i}A_{a}I$ is omitted, the analogue of $\mathbb{M}$ found in [21] is only non-negative definite when a modified dominant energy condition holds - their equation 1.2. However, their condition is very unnatural; for example, it can be seen from the constraint equations that in a purely electrovacuum spacetime with non-zero magnetic field, their modified dominant energy condition never holds.
The closest paper to the present work is [22] - see their section 3.1 in particular. They have the same connection as used here and make similar observations about the decay rates in equation 5.37. However, their main result - their equation 22 - does not include the fourth term in lemma 5.2 because they relied on [20] to get their result. A heuristic argument is given describing corrections to [20], leading to the same conclusions as corollary 5.2.1 (for $S^{2}$ cross-section topology) and equation 5.42 below, but the analytic steps required - including the Dirac equation analysis in appendix A - are not given in [22].
@@
It turns out an explicit locally supersymmetric solution - see section 2.3.1 of [55] or appendix B of [56] - is known to 5D minimal gauged supergravity with $\mathbb{R}\times L(p,1)$ conformal infinity. However, that solution satisfies a different BPS-like equation, namely
@@
Finally, the only other relevant boundary considered in section 4 is the torus. In this case, theorem 4.2 is unchanged by the electromagnetic fields in either theory because $\bar{\varepsilon}_{k}\varepsilon_{k}$ is simply zero for the required $\varepsilon_{k}$.
@@ 6 Acknowledgements
Most of all, I would like to thank my supervisor, James Lucietti, for guidance and many discussions throughout this project. I would also like to thank Jelle Hartong, Hari Kunduri, Harvey Reall and especially Piotr Chrusciel for comments and discussions based on earlier versions of this paper. Finally, I would like to thank the University of Edinburgh’s School of Mathematics for my PhD studentship funding.
@@ Appendix A Inverting the Dirac operator
The key technical requirement is a weighted Poincaré inequality. For the chosen asymptotics, $\Sigma_{t}$ satisfies definition 9.9 of [58] to be a weakly, asymptotically hyperboloidal end [^21] The $x$ in their definition is $\mathrm{e}^{-r}$ here, their $h$ is the pullback of $f_{mn}$ to $\Sigma_{t}$ here and their $\mathcal{N}$ is $S$ here.. Then, one can apply proposition 8.3 of [58] to deduce $\exists\,w\in L_{\mathrm{loc}}^{1}$ such that
@@
for any $\psi\in C_{c}^{\infty}$. Thus,
@@
$\mathfrak{D}$ is a continuous, linear isomorphism between $\mathcal{H}$ and $L^{2}$.
@@
The surjectivity proof follows an index theory argument based on the analysis in [59, 46]. Let $\hat{\nabla}_{a}=D_{a}+\frac{\mathrm{i}}{2}\gamma_{a}$, i.e. don’t include the $\mathcal{A}_{a}$ term. Therefore, $\mathfrak{D}-\hat{\mathfrak{D}}=\gamma^{I}\mathcal{A}_{I}$. If $\gamma^{I}\mathcal{A}_{I}$ were a compact operator, then the index (dimension of kernel minus dimension of cokernel) of $\mathfrak{D}$ and $\hat{\mathfrak{D}}$ would coincide. In particular, if $\hat{\mathfrak{D}}$ was invertible, then it would follow that $\mathrm{index}(\mathfrak{D})=0$. Since $\mathfrak{D}$ has already been shown to have trivial kernel above, it would follow that $\mathfrak{D}$ is invertible.
Even though $\mathcal{A}_{a}$ is just a (sufficiently regular) matrix valued function, it’s not clear in general whether $\gamma^{I}\mathcal{A}_{I}$ is actually compact. This would be the case though if the underlying space, $\Sigma_{t}$, were itself compact - a fact leveraged in [59, 46].
[…]
Since $C_{c}^{\infty}$ is dense in $L^{2}$, for any $\Phi\in L^{2}$, $\exists$ a Cauchy sequence, $\{\phi_{m}\}_{m=0}^{\infty}\subseteq C_{c}^{\infty}$, such that $\lim_{m\to\infty}\phi_{m}=\Phi$ in $L^{2}$. For each $\phi_{m}$ choose $r_{m}$ so that $\mathrm{supp}(\phi_{m})\subseteq\mathrm{int}(\Sigma(r_{m}))$, i.e. choose $r_{m}$ large enough. Then, the choice of boundary conditions on $\partial\Sigma(r_{m})$ doesn’t matter and $\exists\psi_{m}$ such that $\mathfrak{D}(\psi_{m})=\phi_{m}$. Theorem 6.4 in [57, 58] shows $\mathfrak{D}$ has an “elliptic regularity” property on $\Sigma(r_{m})$ meaning $\phi_{m}\in C_{c}^{\infty}$ and the metric’s smoothness imply $\psi_{m}\in C_{c}^{\infty}$ [^23] Theorem 6.4 as stated in [57, 58] only says $\psi_{m}\in H^{1}_{\mathrm{loc}}(\Sigma(r_{m}))$. However, this is because of the very low regularity assumed in [57, 58] for the metric coefficients and the analogue of $\phi_{m}$. If one assumes additional regularity - in this case $g$’s smoothness and $\phi_{m}\in C_{c}^{\infty}$ - then $\psi_{m}$ inherits this additional regularity by the same proof. I’d like to thank Piotr Chrusciel for confirming that this is indeed the case..
Furthermore, $\{\psi_{m}\}_{m=0}^{\infty}\subseteq C_{c}^{\infty}$ is a Cauchy sequence by lemma A.4. By theorem A.6, the limit, $\lim_{m\to\infty}\psi_{m}=\psi\in\mathcal{H}$ satisfies
thereby proving $\mathfrak{D}$ is surjective. ∎
The index theory proof of surjectivity presented works for any regular $\mathcal{A}_{I}$. This is different to the Lichnerowicz identity based proofs which have previously appeared in the literature. These proofs require additional assumptions on $\mathcal{A}_{I}$ which the index theory argument does not. For example, the standard surjectivity argument presented in [57, 58] proceeds as follows.
@@
Suppose $\exists\,\tilde{\mathcal{A}}_{a}$ such that $(\gamma^{I}\mathcal{A}_{I})^{\dagger}=\gamma^{I}\tilde{\mathcal{A}}_{I}$. Define a new connection on spinors, $\widetilde{\nabla}_{a}$, by $\widetilde{\nabla}_{a}=D_{a}-\frac{\mathrm{i}}{2}\gamma_{a}+\tilde{\mathcal{A}}_{a}$ where and a new Dirac operator, $\widetilde{\mathfrak{D}}=\gamma^{I}\widetilde{\nabla}_{I}$. Then, equation A.8 can be re-written as
@@
Since $\psi$ could be any compactly supported spinor, it follows that $\Phi$ is a weak solution to $\widetilde{\mathfrak{D}}(\Phi)=0$. Thus, the surjectivity proof reduces to showing $\widetilde{\mathfrak{D}}$ has trivial kernel. Continuing with the new connection, suppose $\gamma^{IJ}\tilde{\mathcal{A}}_{J}$ is hermitian and
is non-negative definite. Then, following the same steps as theorem 2.6, one finds if $\psi\in C_{c}^{\infty}$, then
@@
As $\widetilde{\mathbb{M}}$ is assumed to be non-negative definite, this is formally identical to the original Lichnerowicz identity. Then, one can proceed analogously to the analysis of $\mathfrak{D}$ itself by defining an inner product analogous to equation A.1 and thereby concluding that $\widetilde{\mathfrak{D}}$ is injective. Hence, $\widetilde{\mathfrak{D}}(\Phi)=0$ implies $\Phi=\theta-\mathfrak{D}(\varphi)=0$ and therefore $\mathfrak{D}$ is surjective.
While this proof is more in the spirit of the rest of the material because it is built on a Lichnerowicz identity, it has the disadvantage that there may be choices of $\mathcal{A}_{a}$ for which no simple $\tilde{\mathcal{A}}_{a}$ exists. Furthermore, even if $\tilde{\mathcal{A}}_{a}$ exists, it may be that $\widetilde{\mathbb{M}}$ is not non-negative definite.
@@
Unlike the proof of theorem A.7 presented above, it appears the Licherowicz identity based proof can only be rectified in a few specific cases where further assumptions are made. The simplest, but most unsatisfying, assumption would be to restrict to electromagnetic fields which have vanishing Poynting vector.